A four year loan of $5000 is repaid by equal annual payments at the end of each year. Compute the annual payment on the basis of an interest rate of 6% p.a. and draw up a loan
schedule, showing the interest component of every payment and the outstanding balance.
1
Expert's answer
2020-10-01T17:27:43-0400

solution


Loan amount "A=5000"

Time "n=4"

Interest "I=6\%"

Payments "P=?"



"A=P * \frac{1-(1+I)^{-n}}{I}"

"5000=P * \frac{1-(1.06)^{-4}}{0.06}"

"5000=P*3.4651"

"P=1442.9575"

the annual payment amount is $1442.96


Each annual payment pays interest on the outstanding loan balance and a portion of the loan balance.


Interest paid is calculated by multiplying the interest rate by the loan balance from the previous period.


"Interest\ paid = Interest (I)*Loan\ Balance"

Principle repaid is the portion of the annual payment that covers part of the loan amount.

"Principle\ repaid= payment\ amount-Interest \ paid"

Loan balance for a given year is the difference between the loan balance from the previous year and the portion of the loan amount repaid during the year


"Loan\ balance_{year\ t} = Loan\ balance_{year\ t-1} - principle\ repaid_{year\ t}"

Note that at the beginning of the loan term, loan balance is equal to the loan amount.


Loan schedule:


year 1:

payment amount = $1442.96

interest paid

"=5000*0.6"

= $300


principle repaid

"=1442.96-300"

= $1142.96


Loan balance

"=5000- 1142.96"

= $3857.04


year 2:

payment amount = $1442.96

interest paid

"=3857.04*0.06"

= $231.42


principle repaid

"=1442.96-231.42"

= $1211.54


Loan balance

"=3857.04-1211.54"

= $ 2645.50


year 3:

payment amount = $1442.96

interest paid

"=2645.50*0.06"

= $158.73


principle repaid

"=1442.96-158.73"

= $1284.23


Loan balance

"=2645.50-1284.23"

= $1361.27


year 4:

payment amount = $1442.96

interest paid

"=1361.27*0.06"

= $81.68


principle repaid

"=1442.96-81.68"

= $1361.28

Loan balance = 0.00